Jeff and Sharon drove for 4 hours with a wind speed of 10 mph. Traveling against the same wind speed, it took them 6 hours to make the return trip. How fast were they traveling in still air?

Answer by stanbon(75887)

Jeff and Sharon drove for 4 hours with a wind speed of 10 mph. Traveling against the same wind speed, it took them 6 hours to make the return trip. How fast were they traveling in still air?
---------------
With-wind DATA:
time = 4 hrs ; rate = x+10 mph ; distance = tr = 4(x+10) miles
----
Against-wind DATA:
time = 6 hrs ; rate = x-10 mph ; distance = tr = 6(x-10) miles
----
Equation:
distance = distance
4(x+10) = 6(x-10)
4x+40 = 6x-60
2x = 100
x = 50 mph (rate in still air)
======
Cheers,
Stan H.